# Improved subsample-and-aggregate via the private modified winsorized mean

### v1

@@ 2.4 Private quantile estimation
A private version of the trimmed mean, $\hat{\mu}_{p}$, requires differentially private estimation of extreme quantiles. We make use of the unbounded private quantile estimator, developed by Durfee, [13] and studied briefly by Ramsay and Diaz-Rodriguez, [31]. We chose the unbounded quantile estimator as it has been shown to estimate extreme quantiles consistently, whereas other private quantile estimators do not [31]. Specifically, when there is no contamination, our mean estimator requires a private analogue of an $O(1/n)$ quantile. It is expected that a private $O(1/n)$th quantile would be consistent for the infimum of the support of $F$, as $n\to\infty$. [31] show that other private quantile estimators do not have this property, however, the unbounded estimator does.
We present the modified estimator given by Ramsay and Diaz-Rodriguez, [31]. Let $V,V_{1},V_{2},\ldots$ be a sequence of independent, standard exponential random variables. Suppose that we wish to estimate the $q$th quantile of the population distribution privately, where the desired quantile is an upper quantile, (i.e., $1/2\leq q\leq 1$). First, we supply the algorithm with the following input parameters: a grid size parameter $\beta>1$, a lower bound on the desired quantile $\ell$, a dataset with empirical distribution $F_{n}$, and two privacy budgets $\varepsilon_{1},\varepsilon_{2}>0$. Given these parameters, the algorithm proceeds as follows:
  1. 1.
    Compute $\hat{q}=q+V/n\varepsilon_{1}$ and set $i=1$.
  2. 2.
    Increase $i$ by one until $F_{n}(\beta^{i}+\ell-1)+V_{i}/n\varepsilon_{2}>\hat{q}$. That is, stop at the first $i$ such that $F_{n}(\beta^{i}+\ell-1)+V_{i}/n\varepsilon_{2}>\hat{q}$.
  3. 3.
    Output $\beta^{i}+\ell-1$.
If $20<q<1/$ 2, we supply an upper bound on the desired quantile, $u$. Then, we negate the dataset, and run the upper-quantile algorithm, with input quantile $1-q$ and $\ell=-u$. This produces a negated estimate of the lower quantile that we negate to get the final quantile estimate. We refer to these quantile estimates by $\tilde{\xi}_{q}\coloneqq\tilde{\xi}_{q,\mathbf{X}_{n}}(\varepsilon_{1},\varepsilon_{2},\ell,u,\beta)$. The algorithm is $(\varepsilon_{1}+\varepsilon_{2})$-differentially private [13].
We show (Lemma D.1) that replacing $\varepsilon_{i}$ with $\sqrt{\rho_{i}}$ and taking $V,V_{1},V_{2},\ldots$ to be standard Gaussian random variables produces $(\rho_{1}+\rho_{2})$-zero-concentrated differentially private quantiles. We denote the estimates generated by the zero-concentrated differentially private version of the algorithm by $\tilde{\xi}_{q}^{\prime}\coloneqq\tilde{\xi}_{q,\mathbf{X}_{n}}^{\prime}(\rho_{1},\rho_{2},\ell,u,\beta)$. As a by-product of our analysis, we prove a simple concentration result concerning both versions of these private quantiles (Lemmas E.1-E.2).
@@ 3 A new differentially private mean estimator
3.1 Definition
We now introduce our private mean estimator, the private modified winsorized mean, drawing inspiration from Lugosi and Mendelson, [28] and Durfee, [13]. Given a trimming proportion $20<p<1/$ 2 and private quantile parameters $\ell,u\in\mathbb{R}$, $\beta>1$, and $\varepsilon_{1},\varepsilon_{2}>0$ (PDP) or $\rho_{1},\rho_{2}>0$ (zCDP), we first compute $\tilde{\xi}_{p}$ and $\tilde{\xi}_{1-p}$ (PDP) or $\tilde{\xi}_{p}^{\prime}$ and $\tilde{\xi}_{1-p}^{\prime}$ (zCDP). Recall that these are the private estimates of the extreme quantiles, based on $\mathbf{Y}_{n}^{\prime}$. The data $\mathbf{X}_{n}^{\prime}$ are then projected onto the interval $[\tilde{\xi}_{p},\tilde{\xi}_{1-p}]$ (PDP) or $[\tilde{\xi}_{p}^{\prime},\tilde{\xi}_{1-p}^{\prime}]$ (zCDP). Lastly, we take the sample mean of the projected observations, and add calibrated noise via the Laplace mechanism (PDP) or the Gaussian mechanism (zCDP). Given $14e^{-n}<\delta<$ 1, our main result is a bound on the deviations of the private modified winsorized mean about $\mu$ that holds with probability $1-\delta$. Under two mild conditions, taking
$p=\zeta\coloneqq\zeta(n,\eta,\delta,\ell,u,\beta)=16\eta+\frac{112}{3}\frac{\log(32(\beta(u-\ell)/(\beta-1)\vee 1)/\delta)}{n},$
for both PDP and zCDP, yields minimax optimal estimates of $\mu$ for the problem stated in Condition 1, where the class of estimators is constrained to those that are differentially private (PDP or zCDP). Here, for two real numbers $a$ and $b$, we write $a\wedge b$ ($a\vee b$) to denote the minimum (maximum) of $a$ and $b$. We formally define the estimators as follows.
Definition 3.1.
Given private quantile parameters $\ell,u\in\mathbb{R}$, $\beta>1$, and $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}>0$ (PDP) or $\rho_{1},\rho_{2},\rho_{3}>0$ (zCDP), the private modified winsorized means are
$\tilde{\mu}=\sum_{i=1}^{n}\frac{1}{n}\phi_{\tilde{\xi}_{\zeta},\tilde{\xi}_{1-\zeta}}(X_{i}^{\prime})+Z_{1}\frac{(\tilde{\xi}_{1-\zeta}-\tilde{\xi}_{\zeta})}{n\varepsilon_{3}}\qquad\text{and}\qquad\tilde{\mu}^{\prime}=\sum_{i=1}^{n}\frac{1}{n}\phi_{\tilde{\xi}_{\zeta}^{\prime},\tilde{\xi}_{1-\zeta}^{\prime}}(X_{i}^{\prime})+Z_{2}\frac{(\tilde{\xi}_{1-\zeta}^{\prime}-\tilde{\xi}_{\zeta}^{\prime})}{n\sqrt{2\rho_{3}}},$
where $Z_{1}$ is a standard Laplace random variable and $Z_{2}$ is a standard Gaussian random variable.
It follows from basic composition (Proposition 2.1) that $\tilde{\mu}$ is $(2\varepsilon_{1}+2\varepsilon_{2}+\varepsilon_{3})$-differentially private and that $\tilde{\mu}^{\prime}$ is $(2\rho_{1}+2\rho_{2}+\rho_{3})$-differentially private. Our main result requires two conditions on $\zeta$, $F$, and the input parameters of the private quantile algorithm $(\ell,u,\beta)$.
@@ 3.1 Definition
It holds that $\ell\leq\xi_{5\zeta/4}$, $u\geq\xi_{1-5\zeta/4}$, and $\zeta<1/2$.
Condition 2 requires that the interval $(\ell,u)$ contains the interval $(\xi_{q/4},\xi_{1-q/4})$, which, colloquially, means “most of the uncontaminated data”. Our theoretical results show that the bounds $\ell\leq\xi_{q/4}$, $u\geq\xi_{1-q/4}$ can be loose, replaced by loose bounds on $\mu$ and $\sigma$, or replaced with a weaker condition, see Remark 3.3 and Remark 3.4.
@@
It holds that
$1<\beta<1+\left[\frac{\xi_{5\zeta/4}-\xi_{3\zeta/4}}{u-\xi_{5\zeta/4}+1}\wedge\frac{\xi_{1-3\zeta/4}-\xi_{1-5\zeta/4}}{\xi_{1-5\zeta/4}-\ell+1}\right]\coloneqq 1+b_{n}.$
@@
Condition 3 says that the grid size used for quantile estimation cannot be too coarse. Combining this condition with our main result yields a range of $\beta$ values that produce minimax optimal estimates, up to logarithmic terms, see Remark 3.2. We now introduce our main theoretical result. For quantities $a,b$, we write $a\lesssim b$ ($a\gtrsim b$) when there is a universal constant $C>0$ such that $a\leq Cb$ ($a\geq Cb$).
Theorem 3.1.
@@
Take $\eta\geq 0$, $14e^{-n}<\delta<$ 1, $\varepsilon_{1},\varepsilon_{2}>112/3$, $\rho_{1},\rho_{2}>112/3$ and $\varepsilon_{3},\rho_{3}>0$. For all $n\geq 1$, $u,\ell\in\mathbb{R}$, and $\beta>1$ such that Conditions 13 hold, with probability at least $1-\delta,$
@@
$|\tilde{\mu}-\mu|\lesssim\sigma\left(\sqrt{\eta}\vee\sqrt{\frac{\log(u-\ell)+\log(\beta/(\beta-1))+\log(4/\delta)}{n}}\vee\frac{1}{\sqrt{n}\varepsilon_{3}}\right),$ (3.1)
and
$|\tilde{\mu}^{\prime}-\mu|\lesssim\sigma\left(\sqrt{\eta}\vee\sqrt{\frac{\log(u-\ell)+\log(\beta/(\beta-1))+\log(4/\delta)}{n}}\vee\frac{1}{\sqrt{n\rho_{3}}}\right).$ (3.2)
@@
The proof of Theorem 3.1 is in Appendix C. To prove Theorem 3.1 we first prove a concentration inequality on the private quantiles. As a result, it suffices to bound the quantity $n^{-1}\sum_{i=1}^{n}\phi_{\xi_{\zeta},\xi_{1-\zeta}}(X_{i}^{\prime})-\mu$. For this task, the proof is similar to the proof of Theorem 1 of Lugosi and Mendelson, [28]. However, their proof contains an error, that is resolved here, see Appendix F.
@@
Interpreting Theorem 3.1, we first observe that, provided that $\log(\beta(u-\ell)/(\beta-1))\lesssim\log(4/\delta)$, and Conditions 2 and 3 are satisfied, the rate given in Theorem 3.1 is minimax optimal, regardless of the level of contamination. Observe that the upper bounds in Theorem 3.1 each have three terms: the contamination cost, the sampling error cost, and the privacy cost, listed in order of appearance. The contamination cost and the sampling error cost, together, match the minimax lower bound in the non-private setting [28]. Furthermore, for the PDP mean, the sampling error cost and the privacy cost match the minimax lower bound in the private setting [23]. Using these results together, the private modified winsorized means are minimax optimal.
To see this, note that when there is no contamination, $\eta=0$, the rates match the established minimax bounds for PDP mean estimation [23]. Neither of the components that remain with $\eta=0$ depend on $\eta$, and as such, their contribution to the rate does not change based on the level of contamination. Moreover, ignoring the privacy error, the rates match the established minimax optimal rates for non-private estimation, at any level of contamination [28]. As a result, the PDP modified winsorized means then has to be minimax optimal for all levels of contamination, provided $\eta$ is such that Conditions 2 and 3 are satisfied. Finally, given that PDP lower bounds imply zCDP lower bounds, we have that the zCDP modified winsorized mean is also minimax optimal for this problem.
Remark 3.2 (On choosing $\beta$).
@@
For minimax optimality, we require $\log(\beta/(\beta-1))\lesssim\log(4/\delta)$. Combining this with Condition 3 yields a recommended range for the grid size: $(1-(\delta/4)^{s})^{-1}\leq 1+(\delta/4)^{s}/(1-(\delta/4)^{s})\leq\beta\leq 1+b_{n}$, for some $s>0$. For example, we can select $\beta=O(1+n^{-s}/(1-n^{-s}))$, provided $\delta<n^{-s}<b_{n}$ for some $s>1$. This ensures that the bound (3.1) holds with probability at least $1-n^{-s}$. Regarding $b_{n}$, bounded support makes for a smaller $b_{n}$, while unbounded support, with heavy tails, makes for a larger $b_{n}$. From the proof of Theorem 3.1, we only need the private quantiles to be in some range of extreme quantiles of the population distribution $(\xi_{\zeta/4},\xi_{4\zeta})$ for the lower quantile and $(\xi_{1-4\zeta},\xi_{1-\zeta/4})$ for the upper quantile. This range is large for distributions with unbounded support, and small for those with bounded support. For instance, for the standard uniform distribution, $b_{n}=O(1/n)$, for the standard exponential distribution, $b_{n}\gtrsim 1/(u-\log n+1)$, and for Student’s $t$-distribution with three degrees of freedom, $b_{n}\gtrsim 1,$ when $u=2\xi_{1-5\zeta/4}$ and $\ell=-u$. Thus, we must take a finer grid when applying the procedure to distributions with bounded support, but a coarse grid is acceptable for distributions with unbounded support.
Remark 3.3 (On choosing $u$ and $\ell$).
@@
Condition 2 requires that $\ell<\xi_{5\zeta/4}<\xi_{1-5\zeta/4}<u$. Then, (3.1) and (3.2) imply that taking $\log(u-\ell)\lesssim\log(1/\delta)$ produces a modified winsorized mean that is minimax optimal. Choosing $\ell,u$ thus amounts to ensuring these two bounds are satisfied. First, observe that, in general, we take $\delta$ to be small, say, $\delta=n^{-k}$ for some $k>0$. The bound $\log(u-\ell)\lesssim\log(1/\delta)$ then requires that $u-\ell$ is at most polynomial in $n$. Now, we need $u,\ell$ such that $\ell<\xi_{5\zeta/4}<\xi_{1-5\zeta/4}<u$. For $F$ with bounded support, this is straightforward. If the support is unbounded, we can specify $u,\ell$ according to the following argument. We have that, by Chebyshev’s inequality, $5\zeta/4\leq\sigma^{2}/(\xi_{1-5\zeta/4}-\mu)^{2}$. This implies that $\xi_{1-5\zeta/4}\lesssim\sigma/\zeta^{1/2}+\mu$. Therefore, Condition 2 is implied by $u\gtrsim\sigma/\zeta^{1/2}\vee\mu$. Then, using $\delta=n^{-k}$, $\sigma/\zeta^{1/2}\lesssim\sigma\sqrt{n/k\log n}.$ Assuming that $\mu,\sigma$ are constant in $n$, we take $u\geq n^{1/2+r},$ $\ell\leq-n^{1/2+r},$ for some $r>0$. For large $n$, this implies Condition 2. If one has instead (potentially very loose) bounds such that $-\mu_{0}<\mu<\mu_{0}$ and $\sigma<\sigma_{0}$, then taking $u\gtrsim\sqrt{n\sigma_{0}}\vee\mu_{0}$ and $\ell\lesssim-(\sqrt{n\sigma_{0}}\vee\mu_{0})$ implies Condition 2, and ensures that $u-\ell$ is at most polynomial in $n$.
Remark 3.4 (Weakening Condition 2).
@@
Condition 2 requires that the interval $(\ell,u)$ contains most of the uncontaminated data. This condition can be relaxed to only require that $u>\xi_{5\zeta/4}$ and $\ell<\xi_{1-5\zeta/4}$, (i.e., $u$ is an upper bound on the lower quantile and $\ell$ is a lower bound on the upper quantile), which is weaker. This will come at the expense of replacing $u-\ell$ in (3.1) and (3.2) with $\log n$. This modification still constitutes a minimax optimal rate, up to logarithmic terms. However, given that the bounds on the data can be very loose, as per (3.1) and (3.2), and the results in Section 4, we leave Condition 2 unchanged.
Remark 3.5 (Subgaussian distributions and more).
@@
As usual, (e.g., as in the work of Lugosi and Mendelson, [28]), we can derive stronger bounds under stronger assumptions. Concretely, we use Chebyshev’s inequality to generate the first and last terms in (3.1) and (3.2). This can be replaced with the tightest tail bound given the assumptions. For instance, if we instead assume the uncontaminated data are subgaussian observations, we use a subgaussian tail bound to get tighter deviation bounds. In that case, the private trimmed mean also achieves optimal rates.
Remark 3.6 (Multivariate case).

### v3

@@ 2.4 Private quantile estimation
A private version of the clipped mean, $\hat{\mu}_{p}$, requires differentially private estimation of extreme quantiles. We make use of the unbounded private quantile estimator, developed by Durfee, (2024). We use the unbounded quantile estimator as it has been shown to estimate extreme quantiles consistently, whereas other private quantile estimators do not (Ramsay and Diaz-Rodriguez,, 2024). Specifically, when there is no contamination, our mean estimator requires a private analogue of an $O(1/n)$ quantile. It is expected that a private $O(1/n)$th quantile would be consistent for the infimum of the support of $F$, as $n\to\infty$. Ramsay and Diaz-Rodriguez, (2024) show that other private quantile estimators do not have this property, however, the unbounded estimator does.
We present the modified estimator given by Ramsay and Diaz-Rodriguez, (2024). Let $V,V_{1},V_{2},\ldots$ be a sequence of independent, standard exponential random variables. Suppose that we wish to estimate the $q$th quantile of the population distribution privately, where the desired quantile is an upper quantile, (i.e., $1/2\leq q\leq 1$). First, we supply the algorithm with: a grid size parameter $\beta>1$, a lower bound on the desired quantile $\ell$, a dataset with empirical distribution $F_{n}$, and two privacy budgets $\varepsilon_{1},\varepsilon_{2}>0$. Given these parameters, the algorithm proceeds as follows: (i) Compute $\hat{q}=q+V/n\varepsilon_{1}$ and set $i=1$. (ii) Increase $i$ by one, stopping at the first $i$ such that $F_{n}(\beta^{i}+\ell-1)+V_{i}/n\varepsilon_{2}>\hat{q}$. (iii) Output $\beta^{i}+\ell-1$. If $20<q<1/$ 2, we supply an upper bound on the desired quantile, $u$. Then, we negate the dataset, and run the upper-quantile algorithm, with input quantile $1-q$ and $\ell=-u$. This produces a negated estimate of the lower quantile that we negate to get the final quantile estimate. We refer to these quantile estimates by $\tilde{\xi}_{q}\coloneqq\tilde{\xi}_{q,\mathbf{X}_{n}}(\varepsilon_{1},\varepsilon_{2},\ell,u,\beta)$. The algorithm is $(\varepsilon_{1}+\varepsilon_{2})$-differentially private (Durfee,, 2024).
We show (Lemma 1) that replacing $\varepsilon_{i}$ with $\sqrt{\rho_{i}}$ and taking $V,V_{1},V_{2},\ldots$ to be standard Gaussian random variables produces $(\rho_{1}+\rho_{2})$-zero-concentrated differentially private quantiles. We denote the estimates generated by the zero-concentrated differentially private version of the algorithm by $\tilde{\xi}_{q}^{\prime}\coloneqq\tilde{\xi}_{q,\mathbf{X}_{n}}^{\prime}(\rho_{1},\rho_{2},\ell,u,\beta)$. As a by-product of our analysis, we prove a simple concentration result concerning both versions of these private quantiles (Lemmas 2-3).
@@ 3 A new differentially private mean estimator
We now introduce our private mean estimator, the PMW mean, drawing inspiration from Lugosi and Mendelson, (2021) and Durfee, (2024). Given a clipping proportion $20<p<1/$ 2 and private quantile parameters $\ell,u\in\mathbb{R}$, $\beta>1$, and $\varepsilon_{1},\varepsilon_{2}>0$ (PDP) or $\rho_{1},\rho_{2}>0$ (zCDP), we first compute $\tilde{\xi}_{p}$ and $\tilde{\xi}_{1-p}$ (PDP) or $\tilde{\xi}_{p}^{\prime}$ and $\tilde{\xi}_{1-p}^{\prime}$ (zCDP). Recall that these are the private estimates of the extreme quantiles, based on $\mathbf{Y}_{n}^{\prime}$. The data $\mathbf{X}_{n}^{\prime}$ are then projected onto the interval $[\tilde{\xi}_{p},\tilde{\xi}_{1-p}]$ (PDP) or $[\tilde{\xi}_{p}^{\prime},\tilde{\xi}_{1-p}^{\prime}]$ (zCDP). Lastly, we take the sample mean of the projected observations, and add calibrated noise via the Laplace mechanism (PDP) or the Gaussian mechanism (zCDP). Given $14e^{-n}<\delta<$ 1, our main result is a bound on the deviations of the PMW mean about $\mu$ that holds with probability $1-\delta$. Here, for two real numbers $a$ and $b$, we write $a\wedge b$ ($a\vee b$) to denote the minimum (maximum) of $a$ and $b$. Under two mild conditions, PMW mean is minimax optimal for the problem stated in Condition 1, where the class of estimators is constrained to those that are differentially private (PDP or zCDP), if:
$p=\zeta=16\eta+\frac{112}{3}\frac{\log(32(\beta(u-\ell)/(\beta-1)\vee 1)/\delta)}{n}.$
For notational simplicity, we hide the dependence of $\zeta$ on other parameters, but formally, $\zeta\coloneqq\zeta(n,\eta,\delta,\ell,u,\beta)$. We formally define the estimators as follows.
Definition 2.
Given private quantile parameters $\ell,u\in\mathbb{R}$, $\beta>1$, and $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}>0$ (PDP) or $\rho_{1},\rho_{2},\rho_{3}>0$ (zCDP), the PDP PMW mean is
$\tilde{\mu}=\sum_{i=1}^{n}\frac{1}{n}\phi_{\tilde{\xi}_{\zeta},\tilde{\xi}_{1-\zeta}}(X_{i}^{\prime})+Z_{1}\frac{(\tilde{\xi}_{1-\zeta}-\tilde{\xi}_{\zeta})}{n\varepsilon_{3}},$
and the zCDP mean $\mu^{\prime}$ is defined analogously, replacing $\tilde{\xi}_{.}$, $\varepsilon_{3}$, and $Z_{1}$ with $\tilde{\xi}_{.}^{\prime}$, $\sqrt{2\rho_{3}}$, and $Z_{2}$. Here $Z_{1}$ is a standard Laplace random variable and $Z_{2}$ is a standard Gaussian random variable.
It follows from basic composition (Proposition 1) that $\tilde{\mu}$ is $(2\varepsilon_{1}+2\varepsilon_{2}+\varepsilon_{3})$-differentially private and that $\tilde{\mu}^{\prime}$ is $(2\rho_{1}+2\rho_{2}+\rho_{3})$-differentially private. Our main result requires two conditions on $\zeta$, $F$, and the input parameters of the private quantile algorithm $(\ell,u,\beta)$.
@@
$\ell\leq\xi_{5\zeta/4}$, $u\geq\xi_{1-5\zeta/4}$, and $\zeta<1/2$.
Condition 2 requires that the interval $(\ell,u)$ contains the interval $(\xi_{5\zeta/4},\xi_{1-5\zeta/4})$, which, colloquially, means “most of the uncontaminated data”. Our theoretical results show that the bounds $\ell\leq\xi_{5\zeta/4}$, $u\geq\xi_{1-5\zeta/4}$ can be loose, replaced by loose bounds on $\mu$ and $\sigma$, or replaced with a weaker condition, see Remark 2 and Remark 3.
@@
It holds that $\beta>1$ and that
$\beta<1+\left[\frac{\xi_{5\zeta/4}-\xi_{3\zeta/4}}{u-\xi_{5\zeta/4}+1}\wedge\frac{\xi_{1-3\zeta/4}-\xi_{1-5\zeta/4}}{\xi_{1-5\zeta/4}-\ell+1}\right]\coloneqq 1+b_{n}.$
@@
Condition 3 says that the grid size used for quantile estimation cannot be too coarse. Combining this condition with our main result yields a range of $\beta$ values that produce minimax optimal estimates, up to logarithmic terms, see Remark 1. We now introduce our main theoretical result. For quantities $a,b$, we write $a\lesssim b$ ($a\gtrsim b$) when there is a universal constant $C>0$ such that $a\leq Cb$ ($a\geq Cb$).
Theorem 1.
@@
Take $\eta\geq 0$, $14e^{-n}<\delta<$ 1, $\varepsilon_{1},\varepsilon_{2},\rho_{1},\rho_{2}>3/56$, and $\varepsilon_{3},\rho_{3}>0$. For all $n\geq 1$, $u,\ell\in\mathbb{R}$, and $\beta>1$ such that Conditions 13 hold, with probability at least $1-\delta,$
@@
$|\tilde{\mu}-\mu|\lesssim\sigma\sqrt{\frac{\log((u-\ell)\times\beta/(\beta-1)\times 4/\delta)}{n}}\\ +\sigma\sqrt{\eta}+\frac{\sigma}{\sqrt{n}\varepsilon_{3}},$ (1)
and $|\tilde{\mu}^{\prime}-\mu|$ is upper bounded (in the sense of $\lesssim$) by the same quantity, with $\varepsilon_{3}$ replaced with $\sqrt{\rho_{3}}$.
@@
The proof of Theorem 1 is in Appendix C. To prove Theorem 1 we first prove a concentration inequality on the private quantiles. After which, it suffices to bound the quantity $n^{-1}\sum_{i=1}^{n}\phi_{\xi_{\zeta},\xi_{1-\zeta}}(X_{i}^{\prime})-\mu$ with high probability. For this task, the proof is similar to the proof of Theorem 1 of Lugosi and Mendelson, (2021). However, their proof contains an error, that is resolved here (Appendix G). A sample complexity analogue of Theorem 1 is given in Appendix A.
@@
Interpreting Theorem 1, we first observe that, provided that $\log(\beta(u-\ell)/(\beta-1))\lesssim\log(4/\delta)$, and Conditions 2 and 3 are satisfied, the rate given in Theorem 1 is minimax optimal, regardless of the level of contamination. The lower bound ensuring minimax optimality can be generated by combining lower bounds from private (Kamath et al.,, 2020) and nonprivate (Lugosi and Mendelson,, 2019) mean estimation.
Remark 1 (On choosing $\beta$).
@@
For minimax optimality, we require $\log(\beta/(\beta-1))\lesssim\log(4/\delta)$. Combining this with Condition 3 yields a recommended range for the grid size: $(1-(\delta/4)^{s})^{-1}\leq 1+(\delta/4)^{s}/(1-(\delta/4)^{s})\leq\beta\leq 1+b_{n}$, for some $s>0$. For example, we can select $\beta=O(1+n^{-s}/(1-n^{-s}))$, provided, for some $s>1$, $\delta<n^{-s}<b_{n}$. This ensures that the bound (1) holds with probability at least $1-n^{-s}$. Regarding $b_{n}$, population densities which have a bounded support have smaller values of $b_{n}$, when compared to ones with an unbounded support. For instance, for the standard uniform distribution, $b_{n}=O(1/n)$, for the standard exponential distribution, $b_{n}\gtrsim 1/(u-\log n+1)$, and for Student’s $t$-distribution with three degrees of freedom, $b_{n}\gtrsim 1,$ when $u=2\xi_{1-5\zeta/4}$ and $\ell=-u$. Thus, we must take a finer grid when applying the procedure to distributions with bounded support, but a coarse grid is acceptable for distributions with unbounded support.
Remark 2 (On choosing $u$ and $\ell$).
@@
Condition 2 requires that $\ell<\xi_{5\zeta/4}<\xi_{1-5\zeta/4}<u$. Then, (1) implies that taking $\log(u-\ell)\lesssim\log(1/\delta)$ produces a PMW mean that is minimax optimal. Choosing $\ell,u$ thus amounts to ensuring these two bounds are satisfied. First, observe that, in general, we take $\delta$ to be small, say, $\delta=n^{-k}$ for some $k>0$. The bound $\log(u-\ell)\lesssim\log(1/\delta)$ then requires that $u-\ell$ is at most polynomial in $n$. Now, we need $u,\ell$ such that $\ell<\xi_{5\zeta/4}<\xi_{1-5\zeta/4}<u$. For $F$ with bounded support, this is straightforward. If the support is unbounded, we can specify $u,\ell$ according to the following argument. We have that, by Chebyshev’s inequality, $5\zeta/4\leq\sigma^{2}/(\xi_{1-5\zeta/4}-\mu)^{2}$. This implies that $\xi_{1-5\zeta/4}\lesssim\sigma/\zeta^{1/2}+\mu$. Therefore, Condition 2 is implied by $u\gtrsim\sigma/\zeta^{1/2}\vee\mu$. Then, using $\delta=n^{-k}$, $\sigma/\zeta^{1/2}\lesssim\sigma\sqrt{n/k\log n}.$ Assuming that $\mu,\sigma$ are constant in $n$, we take $u\geq n^{1/2+r},$ $\ell\leq-n^{1/2+r},$ for some $r>0$. For large $n$, this choice will imply Condition 2. If one has instead (potentially very loose) bounds such that $-\mu_{0}<\mu<\mu_{0}$ and $\sigma<\sigma_{0}$, then taking $u\gtrsim\sqrt{n\sigma_{0}}\vee\mu_{0}$ and $\ell\lesssim-(\sqrt{n\sigma_{0}}\vee\mu_{0})$ implies Condition 2, and ensures that $u-\ell$ is at most polynomial in $n$.
Remark 3 (Weakening Condition 2).
@@
Condition 2 requires that the interval $(\ell,u)$ contains most of the uncontaminated data. This condition can be relaxed to only require that $u>\xi_{5\zeta/4}$ and $\ell<\xi_{1-5\zeta/4}$, (i.e., $u$ is an upper bound on the lower quantile and $\ell$ is a lower bound on the upper quantile). This will come at the expense of replacing $u-\ell$ in (1) with $\log n$. This modification still constitutes a minimax optimal rate, up to logarithmic terms. However, given that the bounds on the data can be very loose, as per (1), and the results in Section 5, we leave Condition 2 unchanged.
Remark 4 (Subgaussian distributions and more).
@@
As usual, (e.g., as in the work of Lugosi and Mendelson, (2021)), we can derive stronger bounds under stronger assumptions. We use Chebyshev’s inequality to generate the first and last terms in (1). This can be replaced with the tightest tail bound given the assumptions. For instance, if we assume the uncontaminated data are subgaussian, we use a subgaussian tail bound to get tighter deviation bounds. In that case, the PMW mean also achieves optimal rates.
Remark 5 (Multivariate case).